The greatest acceleration or deceleration that a train may have is a. The minimum time in which the train may reach from one station to the other separated by a distance d

Grade:11

3 Answers

Vikas TU
4 years ago
Hiii 
The train can get from one station to another station in minimum time having acceleration and deceleration. 
Time = t/2 
Apply the distance formula 
S = 1/2 at ^2 
Find 
Vikas TU
4 years ago
Hii
Find time from the above equation. 
Time t will be equal to 
t = 2√s/a 
Hope this helps
Regard 
Vikas
Cheers 
Arun
4 years ago
Dear student
 
Accn and declrcn each have taken time = t/2
S= 1/2(at/2)*t= at^2/4
t= 2√d/a
 
Hope it helps
 
Regard
Arun